# CECILIA: Gas-Phase Physical Conditions and Multi-Element Chemistry at Cosmic Noon

### v1

@@ 3.1 Electron Densities
The higher ne at Cosmic Noon is potentially linked to the higher SFRs measured in these galaxies (N. A. Reddy et al., 2023b). SFR surface density is proportional to the gas surface density by the Kennicutt-Schmidt law (R. C. Kennicutt, 1998; R. C. Kennicutt & N. J. Evans, 2012), where giant molecular clouds with larger H2 densities could fuel the larger SFR. The massive stars, in turn, ionize the dense molecular cloud and establish a higher initial electron gas density by ne2n. Prior studies have found tentative evidence of a correlation between ne and SFR or SFR surface density, although there are other scenarios that could promote high ne in distant galaxies (see discussion in R. L. Davies et al., 2021; M. W. Topping et al., 2025a). In this respect, CECILIA may provide insight into the ne evolution of typical SFGs at Cosmic Noon (i.e., extending to low stellar mass and SFRs characteristic of the SFMS), although such an analysis is beyond the scope of this paper.
@@ 3.1.1 Density Prioritization
To avoid these systematic uncertainties, we use the multiple ne measurements available to determine an appropriate ne for Te and abundance calculations. We adopt an average of ne[S II] and ne[O II] when both are directly measured, otherwise we simply use ne[S II] when ne[O II] is unavailable. This is similar to the approach proposed by J. E. Méndez-Delgado et al. (2023), although we favor the average of ne[S II] and ne[O II] even when ne[S II] 100 cm-3 because there is evidence of higher density from the [O II] line ratios. The measured and adopted ne in the Te sample are reported in Table 2. Note that a direct ne[S II] is available in all galaxies in the Te sample.
@@ 3.2.1 Temperature Calculations in the CECILIA Galaxies
With the observational setup of CECILIA, it is possible to measure Te from four ions: [O II], [N II], [S III], and [O III]. The temperature-sensitive emission line ratios [N II]5756/[N II]6585, [S III]6314/[S III]9071,9533, and [O III]4364/[O III]4960,5008 are measured purely from the NIRSpec data. Given the nominal wavelength coverage of the NIRSpec G235M grating, it is only possible to measure [O III]4364 in the two CECILIA galaxies at z2.8: C31 and D40. Unfortunately, the [O III] auroral line is not significantly detected in C31, such that D40 is the only galaxy in CECILIA with a measurement of Te[O III] (see N. S. J. Rogers et al., 2024). Combining the NIRSpec [O II]7322,32 auroral line detections with the MOSFIRE [O II]3727,29 nebular emission allows for a constraint on the Te structure of the low-ionization zone. Cutouts of the [S III] and [O II] auroral lines detected in the CECILIA galaxies are provided in Appendix C (Figures 15 and 16). We note that some galaxies with auroral line detections do not have direct Te owing to a lack of coverage for the corresponding nebular lines or H in NIRSpec.
@@
Unlike the ne-sensitive lines, the auroral and nebular CELs required to measure Te have large wavelength separations, necessitating accurate correction for dust attenuation and flux calibration between the different instruments used to acquire the spectra. As discussed in N. S. J. Rogers et al. (2024), the reddening-corrected H I line intensity ratios in D40 deviate from those predicted by Case B recombination, indicating a persistent, wavelength-dependent flux calibration error in the G235M/G395M data or a difference in the shape of the attenuation curve. For reliable Te measurements, a more careful approach is required for measuring the auroral-to-nebular line ratios.
@@
Consistent with the procedures in N. S. J. Rogers et al. (2024) and similar to the method recommended by G. Stasinska (2023) when the shape of the attenuation curve is unknown, we analyze all metal CEL fluxes relative to neighboring H I recombination lines. For CELs and H I lines of similar wavelength, the CEL-to-H I flux ratio is reddening insensitive and eliminates potential systematics owing to wavelength-dependent flux calibration errors. When we require the relative intensities of two metal CELs, we first divide the two CEL-to-H I ratios, then correct the ratio using the emissivities of the two H I lines assuming Case B recombination at Te1.25104 K and the adopted ne of the galaxy. For example, the ratio of [S III]6314/[S III]9533 is measured by first calculating the flux ratios of [S III]6314/H and [S III]9533/P89549. Although these ratios are nearly reddening independent owing to the small wavelength separations between the lines, we apply a reddening correction based on the outlined in §2.7 using the N. A. Reddy et al. (2020) attenuation law. Next, we determine ([S III]6314/H)/([S III]9533/P8) and multiply this ratio by the theoretical H/P8 ratio, ending with [S III]6314/[S III]9533 that is pseudo-reddening corrected. The depth of the CECILIA NIRSpec observations enable this approach, as the faint Paschen lines in the NIR are required to normalize the flux of the [S III] nebular lines.
@@
For Te calculations, we use the PyNeb getTemDen function with the auroral-to-nebular line ratio and the adopted density from §3.1.1. Te uncertainties are calculated using a distribution of 500 auroral-to-nebular line intensity ratios sampled from a normal distribution, where we recalculate Te for each line ratio and take the standard deviation of the ensemble as the uncertainty on Te.
@@ 3.2.2 Simultaneous Temperature Trends
Before the launch of JWST, direct Te measurements in high-z SFGs were limited in number and variety. Despite significant investment in ground-based spectroscopy, [O III]4364 was only observed in a handful of unlensed galaxies at z1.6 (R. L. Sanders et al., 2020; L. Clarke et al., 2023), and most other auroral lines from low-ionization species were undetected (c.f., R. L. Sanders et al., 2023b). The UV O III]1660,66 auroral lines were detected in individual lensed galaxies (e.g., D. P. Stark et al., 2013; B. L. James et al., 2014; D. A. Berg et al., 2018; A. Citro et al., 2024), while stacking techniques probed faint auroral line emission in the UV and optical (C. C. Steidel et al., 2016; R. F. Trainor et al., 2016). The high-quality NIR spectroscopy from JWST has greatly improved our ability to constrain the gas temperature in distant galaxies, although the auroral lines of [O III] remain the most commonly detected in the era of JWST. Even so, the number of high-z galaxies with multiple Te from different ions (N. S. J. Rogers et al., 2024; B. Welch et al., 2024, 2025; T. M. Stanton et al., 2025; T. Morishita et al., 2025; R. L. Sanders et al., 2025; E. Cataldi et al., 2025) is still small relative to the growing archive of Te[O III] measurements.
@@
In addition to the Te trends of the CECILIA galaxies, we plot the direct Te[O II] and Te[S III] measurements from the same z0 EELG sample from Figure 4. We also include the Te measurements in metal-rich H II regions observed as part of the CHemical Abundances of Spirals (CHAOS, D. A. Berg et al., 2015) project. We include the H II regions presented in D. A. Berg et al. (2020), N. S. J. Rogers et al. (2021), and N. S. J. Rogers et al. (2022), where all Te and abundances of the CHAOS regions have been recomputed using the methods described in N. S. J. Rogers et al. (2022). We also plot the D. R. Garnett (1992) photoionization model Te scaling relation as a solid green line. Finally, we include other high-z SFGs with simultaneous Te[S III] and Te[O II] available in the literature, including three SFGs from the MARTA survey reported in E. Cataldi et al. (2025, white triangles) and another eleven galaxies in the AURORA survey from R. L. Sanders et al. (2025, purple triangles). The literature galaxies span a redshift of 1.91 z 4.41, complementary to the CECILIA galaxies.
@@
The CECILIA and literature SFGs in Figure 5 span over 10000 K in Te[O II] and provide insight into the Te structure of high-z galaxies. Figure 5 reveals that the Te[O II]-Te[S III] trends at Cosmic Noon are comparable to the Te observed in local star-forming nebulae. Indeed, five of the six CECILIA galaxies with direct Te[O II] and Te[S III] are within 2 of the trends predicted by the photoionization model Te scaling relation. Similar agreement to local scaling relations was found when comparing the Te[O III]-Te[S III] trends in D40555N. S. J. Rogers et al. (2024) discuss D40’s Te and abundance trends in detail, although our new data reduction indicates Te[O III] is 14300700 K owing to a slightly higher [O III]4363 flux and . This Te is consistent with the original Te[O III] measurement, Te[O III] 13300500 K, and the simultaneous Te[O III] and Te[S III] are in agreement with the D. R. Garnett (1992) scaling relation., the only CECILIA galaxy with a direct Te[O III] measurement (N. S. J. Rogers et al., 2024). Such agreement may suggest that the gas temperatures in the different ionization zones are well described by current photoionization modeling, despite different chemical abundance patterns and ionizing sources in the distant galaxies. As Te is primarily set by the available metals in the ISM, the presence of the abundant O (roughly 37 times more abundant than S in the Sun, M. Asplund et al., 2021) may play a more principal role in shaping the temperature structure in the ISM.
@@
The temperature trends observed in the CECILIA galaxies, local H II regions and EELGs, and those predicted by the photoionization model Te scaling relation suggest that Te[O II] and Te[S III] are comparable from 5000 K up to 15000 K. This behavior is related to the presence of O+ in both the low- and intermediate-ionization zones. Owing to the high excitation energies of the optical CELs, auroral lines are preferentially emitted in the high-Te ISM. This results in a known bias, where the optical auroral lines indicate a higher Te than the volume average temperature (e.g., M. Peimbert, 1967). Since the volume of O+ gas also contains higher ionization ions like S2+, the measurement of Te[O II] may be biased to the higher Te, higher-ionization gas, resulting in Te[O II]-Te[S III] that scatter around the one-to-one line. The Te stratification in the ISM can be significant when comparing direct Te from ions that do not overlap in ionization potential (e.g., see relations in N. S. J. Rogers et al., 2021; J. E. Méndez-Delgado et al., 2023).
@@
When discussing these simultaneous temperatures, it is important to note the dispersion in the direct Te. In local nebulae, Te scaling relations involving Te[O II] generally display large dispersion relative to other common Te scaling relations (e.g., R. C. Kennicutt et al., 2003; D. A. Berg et al., 2020), which is evident in the local nebulae and high-z SFGs plotted in Figure 5. Large differences in Te are even observed when comparing direct Te[O II] to other low-ionization zone Te, such as Te[N II] and Te[S II] (N. S. J. Rogers et al., 2021). Such discrepancies are also observed in the two CECILIA galaxies with direct Te[N II] (BX523 and BX474): in both cases, Te[N II] is determined to be in excess of 1.3104 K (see also A. L. Strom et al., 2021) while Te[O II]104 K and Te[S III]1.1104 K. These two galaxies do not follow the simple assumption that Te[O II]Te[N II] or the Te[N II]-Te[S III] scaling relations available in the literature (e.g., N. S. J. Rogers et al., 2021), either suggesting a more complex Te structure in high-z SFGs or that the large variation in Te[O II] observed in local SFGs is also present at Cosmic Noon. At this time, there are too few high-z SFGs with direct Te[N II] to quantitatively assess the variations in different low-ionization zone temperatures.
@@
Numerous studies have attempted to assess the physical mechanisms influencing the scatter in Te[O II], but no physical condition is clearly correlated with the offset in Te[O II] from the best-fit scaling relation (R. M. Yates et al., 2020; E. Cataldi et al., 2025). This dispersion, therefore, represents a large systematic uncertainty when Te[O II] is unconstrained or required to infer a missing Te, as the functional form of the Te scaling relation cannot capture the empirical scatter in Te[O II] at fixed Te[S III]. Currently, there is insufficient data to derive new empirical Te scaling relations or assess the degree of scatter in Te[O II] at Cosmic Noon. Constraining the shape of these Te scaling relations remains a necessary component for reliable chemical abundance surveys at high-z with future JWST observations.
@@ 3.3 Direct Ionic and Total Abundances
In general, it is possible to measure the ionic abundance of N+, S+, S2+, and Ar2+ using only the NIRSpec observations of the CECILIA galaxies, where measurements of O2+ are possible when z2.32. We measure the abundance of O+ and O2+ (the latter at z2.32) using the MOSFIRE [O II]3727,29 and [O III]5008 emission line fluxes. A measurement of O+ is possible using the auroral [O II] lines, although these lines are relatively faint, are sensitive to variations in Te, and can be produced by dielectronic recombination (R. H. Rubin, 1986; X. W. Liu et al., 2000). This combination leads to potentially large systematic biases in the O+ abundances (M. Rodríguez, 2020; J. E. Méndez-Delgado et al., 2023), so we prioritize the MOSFIRE measurements of the nebular [O II] line intensities relative to H. The exception is when calculating the relative N+/O+ with Te[O II], as the [N II]6585/[O II]7322,32 ratio is measured entirely in NIRSpec, is insensitive to , and is typically measured at higher S/N than [N II]6585/[O II]3727,29.
@@
Ionic abundance calculations are performed using the prioritized density discussed in §3.1.1 and the relevant ionization zone Te. If both Te[N II] and Te[O II] are available, we follow the recommendations of previous chemical abundance surveys and adopt Te[N II] as the low-ionization zone Te (D. A. Berg et al., 2020; N. S. J. Rogers et al., 2021; J. E. Méndez-Delgado et al., 2023). When a direct Te is missing, we use a Te scaling relation to infer the missing ionization zone temperature. For the low- and intermediate- ionization zones, we adopt the empirical scaling relations reported in N. S. J. Rogers et al. (2021), accounting for the intrinsic scatter in Te about the relation when reporting the inferred temperature uncertainty. For the high-ionization zone, we prioritize the D. R. Garnett (1992) scaling relation with Te[S III] when available. This choice is motivated by the agreement between Te[O III]-Te[S III] patterns in local EELGs and the photoionization model relation, with a small intrinsic scatter of 400 K (Rogers et al., in preparation). When Te[S III] is unavailable, we apply the Te[O III]-Te[N II] relation from N. S. J. Rogers et al. (2021) assuming that Te[O II] Te[N II] to infer the high-ionization zone Te.
@@
Total O abundances are determined assuming O/H O+/H+ + O2+/H+, ignoring the potential contribution of O3+ to the total O abundance. In the few instances where the gas-phase O3+/H+ has been constrained in local EELGs, the fractional contribution to the global O/H abundance has been found to be 2% (D. A. Berg et al., 2021; R. J. Rickards Vaught et al., 2025). While [O I] emission is detected in numerous CECILIA galaxies (see Figures 15 and 16), we assume this emission comes from outside the H II regions (either in the PDR or DIG). Finally, we do not account for dust grain depletion of O. This correction can be as large as 0.1 dex (A. Peimbert & M. Peimbert, 2010; M. A. Peña-Guerrero et al., 2012) and is most significant in metal-rich nebulae. This correction may be an important consideration for the most metal-rich CECILIA galaxies, but is likely too large for the metal-poor, highly-ionized ISM observed in the majority of the targets. Therefore, we choose to adopt a consistent approach for O/H calculation and forgo correction for dust depletion.
@@
O is the only element where the most abundant ionization states are directly observable in the rest-optical/NIR. For N, S, and Ar abundances, we require Ionization Correction Factors (ICFs) to correct for the abundance of unobserved ionization states. Empirically-derived ICFs, while available for some elements (e.g., O. L. Dors et al., 2013; C. Esteban et al., 2015), are relatively uncommon owing to the observational challenges associated with detecting all relevant metal ionization states. A simple approach to ICFs is to use ions of similar ionization potential, where it is assumed each ion comprises a similar ionic fraction in the ISM. A notable example that is adopted in the present work is that of N+ and O+, where the relative N/O abundance is often assumed to be N/O N+/O+ and ICF(N) O/O+ (M. Peimbert & R. Costero, 1969). This ICF has been widely adopted for N abundance calculations from the optical [N II] lines in local (L. van Zee & M. P. Haynes, 2006; K. V. Croxall et al., 2015; D. A. Berg et al., 2020) and high-z SFGs (K. Z. Arellano-Córdova et al., 2025; T. Morishita et al., 2025; B. Welch et al., 2025) and is generally valid to within 10% of the true N/O abundance (A. Nava et al., 2006; A. Amayo et al., 2021).
@@
For total S/H and Ar/H, we require ICFs derived with photoionization model predictions. As discussed in N. S. J. Rogers et al. (2024) and in many other chemical abundance studies (see recent works by K. Z. Arellano-Córdova et al., 2024; C. Esteban et al., 2025), the choice of ICF represents a large source of systematic uncertainty when deriving total abundances. This problem is exacerbated at high-z, where typical SFGs are characterized by harder ionizing spectra and distinct chemical abundance patterns (C. C. Steidel et al., 2014; A. E. Shapley et al., 2015, 2024; A. L. Strom et al., 2017, 2018; F. Cullen et al., 2019). The harder ionizing spectra will produce emission line ratios that deviate from those observed in local SFGs; as such, ICFs calibrated on photoionization models selected to reproduce the z 0 BPT locus (e.g., A. Amayo et al., 2021) are inappropriate to apply at high-z. The derivation of new ICFs from a grid photoionization models tailored for the chemical and ionization trends of high-z SFGs is necessary, but beyond the scope of the present work. Instead, we adopt the S and Ar ICFs of Y. I. Izotov et al. (2006), which have been utilized for S and Ar abundances in other high-z SFGs (e.g., T. M. Stanton et al., 2025).
@@
ICF(S) accounts for missing S3+ in the highly-ionized ISM, and we choose a ICF(Ar) that corrects for unobserved Ar+ and Ar3+. This is required because the typical CECILIA galaxies do not have coverage of the optical [Ar IV] emission lines (for tentative detection in D40, see N. S. J. Rogers et al., 2024). The Y. I. Izotov et al. (2006) ICFs are polynomial functions of O+/O, with different calibrations for metal-poor and metal-rich nebulae. None of the CECILIA galaxies are described by 12+log(O/H) 7.2, and we adopt the high-Z calibrations for galaxies with 12+log(O/H) 8.2. For galaxies with 7.6 12+log(O/H) 8.2, we follow the recommendations of Y. I. Izotov et al. (2006) and linearly interpolate between the intermediate- and high-metallicity ICFs. Similar to the methods of N. S. J. Rogers et al. (2024), we assume a 10% uncertainty on the ICF to partially account for the systematic uncertainty related to the choice of ICF. We report the ionic and total abundances for the CECILIA galaxies in Table 2.

### v2

@@ 3.1 Electron Densities
The higher ne at Cosmic Noon is potentially linked to the higher SFRs measured in these galaxies (reddy2023mosdefne). SFR surface density is proportional to the gas surface density by the Kennicutt-Schmidt law (kennicutt1998; kennicutt2012), where giant molecular clouds with larger H2 densities could fuel the larger SFR. The massive stars, in turn, ionize the dense molecular cloud and establish a higher initial electron gas density by ne2n. Prior studies have found tentative evidence of a correlation between ne and SFR or SFR surface density, although there are other scenarios that could promote high ne in distant galaxies (see discussion in davies2021; topping2025). In this respect, CECILIA may provide insight into the ne evolution of typical SFGs at Cosmic Noon (i.e., extending to low stellar mass and SFRs characteristic of the SFMS), although such an analysis is beyond the scope of this paper.
@@ 3.1.1 Density Prioritization
To avoid these systematic uncertainties, we use the multiple ne measurements available to determine an appropriate ne for Te and abundance calculations. We adopt an average of ne[S II] and ne[O II] when both are directly measured, otherwise we simply use ne[S II] when ne[O II] is unavailable. This is similar to the approach proposed by mendez-delgado2023, although we favor the average of ne[S II] and ne[O II] even when ne[S II] 100 cm-3 because there is evidence of higher density from the [O II] line ratios. The measured and adopted ne in the Te sample are reported in Table 2. Note that a direct ne[S II] is available in all galaxies in the Te sample.
@@ 3.2.1 Temperature Calculations in the CECILIA Galaxies
With the observational setup of CECILIA, it is possible to measure Te from four ions: [O II], [N II], [S III], and [O III]. The temperature-sensitive emission line ratios [N II]5756/[N II]6585, [S III]6314/[S III]9071,9533, and [O III]4364/[O III]4960,5008 are measured purely from the NIRSpec data. Given the nominal wavelength coverage of the NIRSpec G235M grating, it is only possible to measure [O III]4364 in the two CECILIA galaxies at z2.8: C31 and D40. Unfortunately, the [O III] auroral line is not significantly detected in C31, such that D40 is the only galaxy in CECILIA with a measurement of Te[O III] (see rogers2024). Combining the NIRSpec [O II]7322,32 auroral line detections with the MOSFIRE [O II]3727,29 nebular emission allows for a constraint on the Te structure of the low-ionization zone. Cutouts of the [S III] and [O II] auroral lines detected in the CECILIA galaxies are provided in Appendix C (Figures 15 and 16). We note that some galaxies with auroral line detections do not have direct Te owing to a lack of coverage for the corresponding nebular lines or H in NIRSpec.
@@
Unlike the ne-sensitive lines, the auroral and nebular CELs required to measure Te have large wavelength separations, necessitating accurate correction for dust attenuation and flux calibration between the different instruments used to acquire the spectra. As discussed in rogers2024, the reddening-corrected H I line intensity ratios in D40 deviate from those predicted by Case B recombination, indicating a persistent, wavelength-dependent flux calibration error in the G235M/G395M data or a difference in the shape of the attenuation curve. For reliable Te measurements, a more careful approach is required for measuring the auroral-to-nebular line ratios.
@@
Consistent with the procedures in rogers2024 and similar to the method recommended by stasinska2023 when the shape of the attenuation curve is unknown, we analyze all metal CEL fluxes relative to neighboring H I recombination lines. For CELs and H I lines of similar wavelength, the CEL-to-H I flux ratio is reddening insensitive and eliminates potential systematics owing to wavelength-dependent flux calibration errors. When we require the relative intensities of two metal CELs, we first divide the two CEL-to-H I ratios, then correct the ratio using the emissivities of the two H I lines assuming Case B recombination at Te1.25104 K and the adopted ne of the galaxy. For example, the ratio of [S III]6314/[S III]9533 is measured by first calculating the flux ratios of [S III]6314/H and [S III]9533/P89549. Although these ratios are nearly reddening independent owing to the small wavelength separations between the lines, we apply a reddening correction based on the outlined in §2.7 using the reddy2020 attenuation law. Next, we determine ([S III]6314/H)/([S III]9533/P8) and multiply this ratio by the theoretical H/P8 ratio, ending with [S III]6314/[S III]9533 that is pseudo-reddening corrected. The depth of the CECILIA NIRSpec observations enable this approach, as the faint Paschen lines in the NIR are required to normalize the flux of the [S III] nebular lines.
@@
For Te calculations, we use the PyNeb getTemDen function with the auroral-to-nebular line ratio and the adopted density from §3.1.1. Te uncertainties are calculated using a distribution of 500 auroral-to-nebular line intensity ratios sampled from a normal distribution, where we recalculate Te for each line ratio and take the standard deviation of the ensemble as the uncertainty on Te. The PyNeb getCrossTemDen function can be used to simultaneously solve for ne and Te using line ratios from ions in the same ionization zone. For galaxies where ne[O II] and Te[O II] can be simultaneously constrained, the [O II] lines produce the same Te and ne determined from the above procedures except for two galaxies. The galaxies BX461 and BX336 have ne[O II] less than the adopted ne from §3.1.1, which produces the difference between the Te[O II] measured from the adopted and getCrossTemDen approaches. We choose to maintain a consistent approach and adopt the average density discussed in §3.1.1 for Te calculations, which leverages the deeper, higher S/N NIRSpec observations to constrain the ISM density when ne[O II] is unavailable.
@@ 3.2.2 Simultaneous Temperature Trends
Before the launch of JWST, direct Te measurements in high-z SFGs were limited in number and variety. Despite significant investment in ground-based spectroscopy, [O III]4364 was only observed in a handful of unlensed galaxies at z1.6 (sanders2020; clarke2023), and most other auroral lines from low-ionization species were undetected (c.f., sanders2023o2). The UV O III]1660,66 auroral lines were detected in individual lensed galaxies (e.g., stark2013; james2014; berg2018; citro2024), while stacking techniques probed faint auroral line emission in the UV and optical (steidel2016; trainor2016). The high-quality NIR spectroscopy from JWST has greatly improved our ability to constrain the gas temperature in distant galaxies, although the auroral lines of [O III] remain the most commonly detected in the era of JWST. Even so, the number of high-z galaxies with multiple Te from different ions (rogers2024; welch2024; welch2025; chakraborty2025; stanton2025; morishita2025; sanders2025; cataldi2025) is still small relative to the growing archive of Te[O III] measurements.
@@
In addition to the Te trends of the CECILIA galaxies, we plot the direct Te[O II] and Te[S III] measurements from the same z0 EELG sample from Figure 4. We also include the Te measurements in metal-rich H II regions observed as part of the CHemical Abundances of Spirals (CHAOS, berg2015) project. We include the H II regions presented in berg2020, rogers2021, and rogers2022, where all Te and abundances of the CHAOS regions have been recomputed using the methods described in rogers2022. We also plot the garnett1992 photoionization model Te scaling relation as a solid green line. Finally, we include other high-z SFGs with simultaneous Te[S III] and Te[O II] available in the literature, including three SFGs from the MARTA survey reported in cataldi2025 and another eleven galaxies in the AURORA survey from sanders2025. The literature galaxies span a redshift of 1.91 z 4.41, complementary to the CECILIA galaxies.
@@
The CECILIA and literature SFGs in Figure 5 span over 10000 K in Te[O II] and provide insight into the Te structure of high-z galaxies. Figure 5 reveals that the Te[O II]-Te[S III] trends at Cosmic Noon are comparable to the Te observed in local star-forming nebulae. Indeed, five of the six CECILIA galaxies with direct Te[O II] and Te[S III] are within 2 of the trends predicted by the photoionization model Te scaling relation. Similar agreement to local scaling relations was found when comparing the Te[O III]-Te[S III] trends in D40555rogers2024 discuss D40’s Te and abundance trends in detail, although our new data reduction indicates Te[O III] is 14300700 K owing to a slightly higher [O III]4363 flux and . This Te is consistent with the original Te[O III] measurement, Te[O III] 13300500 K, and the simultaneous Te[O III] and Te[S III] are in agreement with the garnett1992 scaling relation., the only CECILIA galaxy with a direct Te[O III] measurement (rogers2024). Such agreement may suggest that the gas temperatures in the different ionization zones are well described by current photoionization modeling, despite different chemical abundance patterns and ionizing sources in the distant galaxies. As Te is primarily set by the available metals in the ISM, the presence of the abundant O (roughly 37 times more abundant than S in the Sun, asplund2021) may play a more principal role in shaping the temperature structure in the ISM.
@@
The temperature trends observed in the CECILIA galaxies, local H II regions and EELGs, and those predicted by the photoionization model Te scaling relation suggest that Te[O II] and Te[S III] are comparable from 5000 K up to 15000 K. This behavior is related to the presence of O+ in both the low- and intermediate-ionization zones. Owing to the high excitation energies of the optical CELs, auroral lines are preferentially emitted in the high-Te ISM. This results in a known bias, where the optical auroral lines indicate a higher Te than the volume average temperature (e.g., peimbert1967). Since the volume of O+ gas also contains higher ionization ions like S2+, the measurement of Te[O II] may be biased to the higher Te, higher-ionization gas, resulting in Te[O II]-Te[S III] that scatter around the one-to-one line. The Te stratification in the ISM can be significant when comparing direct Te from ions that do not overlap in IP (e.g., see relations in rogers2021; mendez-delgado2023).
@@
When discussing these simultaneous temperatures, it is important to note the dispersion in the direct Te. In local nebulae, Te scaling relations involving Te[O II] generally display large dispersion relative to other common Te scaling relations (e.g., kennicutt2003; berg2020), which is evident in the local nebulae and high-z SFGs plotted in Figure 5. Large differences in Te are even observed when comparing direct Te[O II] to other low-ionization zone Te, such as Te[N II] and Te[S II] (rogers2021). Such discrepancies are also observed in the two CECILIA galaxies with direct Te[N II] (BX523 and BX474): in both cases, Te[N II] is determined to be in excess of 1.3104 K (see also strom2023) while Te[O II]104 K and Te[S III]1.1104 K. These two galaxies do not follow the simple assumption that Te[O II]Te[N II] or the Te[N II]-Te[S III] scaling relations available in the literature (e.g., rogers2021), either suggesting a more complex Te structure in high-z SFGs or that the large variation in Te[O II] observed in local SFGs is also present at Cosmic Noon. At this time, there are too few high-z SFGs with direct Te[N II] to quantitatively evaluate the variations in different low-ionization zone temperatures.
@@
Numerous studies have attempted to assess the physical mechanisms influencing the scatter in Te[O II], but no physical condition is clearly correlated with the offset in Te[O II] from the best-fit scaling relation (yates2020; cataldi2025). This dispersion, therefore, represents a large systematic uncertainty when Te[O II] is unconstrained or required to infer a missing Te, as the functional form of the Te scaling relation cannot capture the empirical scatter in Te[O II] at fixed Te[S III]. Currently, there is insufficient data to derive new empirical Te scaling relations or assess the degree of scatter in Te[O II] at Cosmic Noon. Constraining the shape of these Te scaling relations remains a necessary component for reliable chemical abundance surveys at high-z with future JWST observations.
@@ 3.3 Direct Ionic and Total Abundances
In general, it is possible to measure the ionic abundance of N+, S+, S2+, and Ar2+ using only the NIRSpec observations of the CECILIA galaxies, where measurements of O2+ are possible when z2.32. We measure the abundance of O+ and O2+ (the latter at z2.32) using the MOSFIRE [O II]3727,29 and [O III]5008 emission line fluxes. A measurement of O+ is possible using the auroral [O II] lines, although these lines are relatively faint, are sensitive to variations in Te, and can be produced by dielectronic recombination (rubin1986; liu2000). This combination leads to potentially large systematic biases in the O+ abundances (rodriguez2020; mendez-delgado2023), so we prioritize the MOSFIRE measurements of the nebular [O II] line intensities relative to H. The exception is when calculating the relative N+/O+ with Te[O II], as the [N II]6585/[O II]7322,32 ratio is measured entirely in NIRSpec, is insensitive to , and is typically measured at higher S/N than [N II]6585/[O II]3727,29.
@@
Ionic abundance calculations are performed using the prioritized density discussed in §3.1.1 and the relevant ionization zone Te. If both Te[N II] and Te[O II] are available, we follow the recommendations of previous chemical abundance surveys and adopt Te[N II] as the low-ionization zone Te (berg2020; rogers2021; mendez-delgado2023). When a direct Te is missing, we use a Te scaling relation to infer the missing ionization zone temperature. For the low- and intermediate- ionization zones, we adopt the empirical scaling relations reported in rogers2021, accounting for the intrinsic scatter in Te about the relation when reporting the inferred temperature uncertainty. For the high-ionization zone, we prioritize the garnett1992 scaling relation with Te[S III] when available. This choice is motivated by the agreement between Te[O III]-Te[S III] patterns in local EELGs and the photoionization model relation, with a small intrinsic scatter of 400 K (Rogers et al., in preparation). When Te[S III] is unavailable, we apply the Te[O III]-Te[N II] relation from rogers2021 assuming that Te[O II] Te[N II] to infer the high-ionization zone Te.
@@
Total O abundances are determined assuming O/H O+/H+ + O2+/H+, ignoring the potential contribution of O3+ to the total O abundance. In the few instances where the gas-phase O3+/H+ has been constrained in local EELGs, the fractional contribution to the global O/H abundance has been found to be 2% (berg2021; rickardsvaught2025). While [O I] emission is detected in numerous CECILIA galaxies (see Figures 15 and 16), we assume this emission comes from outside the H II regions (either in the PDR or DIG). Finally, we do not account for dust grain depletion of O. This correction can be as large as 0.1 dex (peimbert2010; pena2012) and is most significant in metal-rich nebulae. This correction may be an important consideration for the most metal-rich CECILIA galaxies, but is likely too large for the metal-poor, highly-ionized ISM observed in the majority of the targets. Therefore, we choose to adopt a consistent approach for O/H calculation and forgo correction for dust depletion.
@@
O is the only element where the most abundant ionization states are directly observable in the rest-optical/NIR. For N, S, and Ar abundances, we require Ionization Correction Factors (ICFs) to correct for the abundance of unobserved ionization states. Empirically-derived ICFs, while available for some elements (e.g., dors2013; esteban2015), are relatively uncommon owing to the observational challenges associated with detecting all relevant metal ionization states. A simple approach to ICFs is to use ions of similar IP, where it is assumed each ion comprises a similar ionic fraction in the ISM. A notable example that is adopted in the present work is that of N+ and O+, where the relative N/O abundance is often assumed to be N/O N+/O+ and ICF(N) O/O+ (peimbert1969). This ICF has been widely adopted for N abundance calculations from the optical [N II] lines in local (vanZee2006; croxall2015; berg2020) and high-z SFGs (arellano-cordova2025; morishita2025; welch2025) and is generally valid to within 10% of the true N/O abundance (nava2006; amayo2021).
@@
For total S/H and Ar/H, we require ICFs derived with photoionization model predictions. As discussed in rogers2024 and in many other chemical abundance studies (see recent works by arellano-cordova2024a; esteban2025), the choice of ICF represents a large source of systematic uncertainty when deriving total abundances. This problem is exacerbated at high-z, where typical SFGs are characterized by harder ionizing spectra and distinct chemical abundance patterns (steidel2014; shapley2015; shapley2024; strom2017; strom2018; cullen2019). The harder ionizing spectra will produce emission line ratios that deviate from those observed in local SFGs; as such, ICFs calibrated on photoionization models selected to reproduce the z 0 BPT locus (e.g., amayo2021) are inappropriate to apply at high-z. The derivation of new ICFs from a grid photoionization models tailored for the chemical and ionization trends of high-z SFGs is necessary, but beyond the scope of the present work. Instead, we adopt the S and Ar ICFs of izotov2006, which have been utilized for S and Ar abundances in other high-z SFGs (e.g., stanton2025).
@@
ICF(S) accounts for missing S3+ in the highly-ionized ISM, and we choose a ICF(Ar) that corrects for unobserved Ar+ and Ar3+. This is required because the typical CECILIA galaxies do not have coverage of the optical [Ar IV] emission lines (for tentative detection in D40, see rogers2024). The izotov2006 ICFs are polynomial functions of O+/O, with different calibrations for metal-poor and metal-rich nebulae. None of the CECILIA galaxies are described by 12+log(O/H) 7.2, and we adopt the high-Z calibrations for galaxies with 12+log(O/H) 8.2. For galaxies with 7.6 12+log(O/H) 8.2, we follow the recommendations of izotov2006 and linearly interpolate between the intermediate- and high-metallicity ICFs. Similar to the methods of rogers2024, we assume a 10% uncertainty on the ICF to partially account for the systematic uncertainty related to the choice of ICF. We report the ionic and total abundances for the CECILIA galaxies in Table 2.